I really need help to make sure I did this right. I have to show my work, and I think I did it right. (I hope anyway) Thank you in advance!
Question: If you flipped a coin three times, what would the probability be that you would get heads on all three times. (show work)
My Answer:
P(H)= number of outcomes that are heads = 1
---------------------------------------------
total number of possible outcomes = 3

Answer by Nate(3500)

Number of Outcomes = 1
--------------------------------------
Number of Possibilities = 2^3 = 8
Chance = 1/8 or 12.5%
Possibilites:
HHH*
HHT
HTH
THH
HTT
THT
TTH
TTT